# Flow Matching with Gaussian Process Priors for Probabilistic Time Series Forecasting

### v1

@@ A.5 Conditional Priors
$\mathbf{x}_{i}^{\mathrm{norm}}=\frac{\mathbf{x}_{i}-\mu_{p}}{\sigma_{p}},$ (16)
where $\mu_{p}$ is the mean and $\sigma_{p}$ is the standard deviation computed over segments of length corresponding to the periodicity $p$ of the time series. This normalization ensures that each segment of the data, organized by its periodic frequency, has a mean of zero and a standard deviation of one, aligning the scales across different frequencies within the dataset.
A.6 Runtime
@@
In the unconditional model, we observed a neglectable computational overhead when using optimal transport maps during training. For evaluation, however, it takes approximately three times longer than the standard conditional version for evaluation. We attribute this increase to two factors: the use of a longer context window (see Sec. A.2) and the need to differentiate through the ODE integration process.
A.7 Additional Forecasts
@@
A.8 Algorithms
A.8.1 Training Algorithms for TSFlow
@@
Algorithm 1 Unconditional Training of TSFlow
1:Input: Prior distribution $q_{0}$, data distribution $q_{1}$, noise level $\sigma_{\mathrm{min}}$, network $u_{\bm{\theta}}$
2:for $\mathrm{iteration}=1,\dots$ do
5:     $t\sim\mathcal{U}(0,1)$    $\triangleright$ sample random time step $t$
7:     $\mathbf{x}_{t}\sim\mathcal{N}(\bm{\mu}_{t},\sigma_{\mathrm{min}}^{2}\mathbf{I})$triangleright sample from $p_{t}(\cdot\mid\mathbf{x}_{0},\mathbf{y})$
8:     $\mathcal{L}({\bm{\theta}})\leftarrow\|u_{\bm{\theta}}(\mathbf{x}_{t},t)-(\mathbf{y}-\mathbf{x}_{0})\|^{2}$triangleright regress conditional vector field
10:end for
11:Return: $u_{\bm{\theta}}$
@@
1:Input: Conditional prior distribution $q_{0}$, data distribution $q_{1}$, noise level $\sigma_{\mathrm{min}}$, network $u_{\bm{\theta}}$
@@
6:     $\mathbf{x}_{t}\sim\mathcal{N}(\bm{\mu}_{t},\sigma_{\mathrm{min}}^{2}\mathbf{I})$triangleright sample from $p_{t}(\cdot\mid\mathbf{x}_{0},\mathbf{y})$
7:     $\mathcal{L}({\bm{\theta}})\leftarrow\|u_{\bm{\theta}}(\mathbf{x}_{t},\mathbf{y}^{p},t)-(\mathbf{y}-\mathbf{x}_{0})\|^{2}$triangleright regress conditional vector field
@@
A.9 Evaluation Algorithms for TSFlow
We provide a formal description of the conditional prior sampling (see Alg. 3) algorithm in Alg. 3.
@@ A.9 Evaluation Algorithms for TSFlow
4:     $\mathbf{x}_{0}^{(i+1)}=\mathbf{x}_{0}^{(i)}-\tau\nabla_{\mathbf{x}_{0}}\log q_{0}(\mathbf{x}_{0}^{(i)}\mid\mathbf{y}^{p})+\sqrt{2\tau}\xi_{i}$triangleright update $\mathbf{x}_{0}$
@@ Appendix B Additional Results
B.1 Ablation
B.1.1 Conditional Model
@@
B.1.2 Conditional Prior Sampling
@@
0 $\mathbf{0.049}$±$\pm 0.001$ $0.012$±$\pm 0.001$ $0.308$±$\pm 0.002$ $\mathbf{0.029}$±$\pm 0.001$ $\mathbf{0.371}$±$\pm 0.006$ $0.096$±$\pm 0.000$ $0.154$±$\pm 0.001$ $0.290$±$\pm 0.004$
@@
B.2 Gaussian Process Regression
We report the forecasting results of the Gaussian process regression in Tab. 9.
@@ B.2 Gaussian Process Regression
Kernel Electricity Exchange KDDCup M4(H) Solar Traffic UberTLC Wiki2000
OU $\mathbf{0.053}$ $\mathbf{0.011}$ $\mathbf{0.364}$ $\mathbf{0.032}$ $\underline{0.346}$ $\mathbf{0.154}$ $\mathbf{0.205}$ $\mathbf{0.331}$
SE $\underline{0.055}$ $\underline{0.013}$ $\underline{0.395}$ $\underline{0.037}$ $\mathbf{0.345}$ $\underline{0.163}$ $\underline{0.230}$ $\underline{0.350}$
PE $0.066$ $0.017$ $0.423$ $0.047$ $0.456$ $0.200$ $0.284$ $0.409$
@@
Table 9: Forecasting results (CRPS) of our Gaussian process regression on eight real-world datasets. Best scores in bold, second best underlined.
The results show that the Ornstein-Uhlenbeck kernel yields the best forecasts on 7 out of 8 datasets.
@@ Appendix C Guided Generation
To enable guided generation, we introduce slightly modified conditional probability paths defined as:
$p_{t}(\mathbf{x}_{t}\mid\mathbf{z})=\mathcal{N}\left(\mathbf{x}_{t};(1-t)\mathbf{x}_{0}+t\mathbf{x}_{1},\left((1-t)\sigma_{\mathrm{max}}^{2}+t\sigma_{\mathrm{min}}^{2}\right)\mathbf{I}\right),$ (17)
where $\sigma_{\mathrm{max}}^{2}$ and $\sigma_{\mathrm{min}}^{2}$ are constant variances with $\sigma_{\mathrm{max}}^{2}>\sigma_{\mathrm{min}}^{2}$. Note that when $\sigma_{\mathrm{max}}^{2}=\sigma_{\mathrm{min}}^{2}$, we recover the same probability path as in Sec. 2.2. However, we require $\sigma_{\mathrm{max}}^{2}>\sigma_{\mathrm{min}}^{2}$ to obtain a score function in our vector field.
@@
$u_{t}(\mathbf{x}_{t}\mid\mathbf{z})=\mathbf{x}_{1}-\mathbf{x}_{0}+\frac{\sigma_{\mathrm{min}}^{2}-\sigma_{\mathrm{max}}^{2}}{(1-t)\sigma_{\mathrm{max}}^{2}+t\sigma_{\mathrm{min}}^{2}}\left(\mathbf{x}_{t}-(1-t)\mathbf{x}_{0}-t\mathbf{x}_{1}\right).$ (18)
@@
$u_{t}(\mathbf{x}_{t}\mid\mathbf{z})=\mathbf{x}_{1}-\mathbf{x}_{0}-(\sigma_{\mathrm{min}}^{2}-\sigma_{\mathrm{max}}^{2})\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}\mid\mathbf{z})$ (19)
@@
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}\mid\mathbf{z})=-\frac{\mathbf{x}_{t}-(1-t)\mathbf{x}_{0}-t\mathbf{x}_{1}}{(1-t)\sigma_{\mathrm{max}}^{2}+t\sigma_{\mathrm{min}}^{2}}.$ (20)
@@
$=$ $\int\frac{(\mathbf{x}_{1}-\mathbf{x}_{0}-(\sigma_{\mathrm{min}}^{2}-\sigma_{\mathrm{max}}^{2})\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}\mid\mathbf{z}))p_{t}(\mathbf{x}_{t}\mid\mathbf{z})q(\mathbf{z})}{p_{t}(\mathbf{x}_{t})}\mathrm{d}\mathbf{z}$ (22)
$=$ $\int\frac{(\mathbf{x}_{1}-\mathbf{x}_{0})p_{t}(\mathbf{x}_{t}\mid\mathbf{z})q(\mathbf{z})}{p_{t}(\mathbf{x}_{t})}\mathrm{d}\mathbf{z}-(\sigma_{\mathrm{min}}^{2}-\sigma_{\mathrm{max}}^{2})\int\frac{\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}\mid\mathbf{z})p_{t}(\mathbf{x}_{t}\mid\mathbf{z})q(\mathbf{z})}{p_{t}(\mathbf{x}_{t})}\mathrm{d}\mathbf{z}$ (23)
$=$ $v_{t}(\mathbf{x}_{t})-(\sigma_{\mathrm{min}}^{2}-\sigma_{\mathrm{max}}^{2})\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}),$ (24)
@@
$\tilde{u}_{t}(\mathbf{x}_{t})=\underbrace{v_{t}(\mathbf{x}_{t})-(\sigma_{\mathrm{min}}^{2}-\sigma_{\mathrm{max}}^{2})\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})}_{=u_{t}(x_{t})}-\underbrace{\lambda(\sigma_{\mathrm{min}}^{2}-\sigma_{\mathrm{max}}^{2})}_{:=s}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}^{p}\mid\mathbf{x}_{t}),$ (26)
where $\lambda$ and $s$ are scaling parameters introduced to control the strength of the conditioning. We can summarize this as:
$\tilde{u}_{t}(\mathbf{x}_{t})=u_{t}(\mathbf{x}_{t})-s\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}^{p}\mid\mathbf{x}_{t}).$ (27)
Thus, to sample from the conditional generative model $p_{t}(\mathbf{x}_{t}\mid\mathbf{y}^{p})$, we use the modified vector field $\tilde{u}_{t}(\mathbf{x}_{t})$ from Eq. 27 with our trained vector field $u_{\bm{\theta}}(t,\mathbf{x}_{t})$.

### v2

@@ A.7 Conditional Priors
$\mathbf{x}_{i}^{\mathrm{norm}}=\mathbf{x}_{i}-\mu_{p},$ (18)
where $\mu_{p}$ is the mean computed over segments of length corresponding to the periodicity $p$ of the time series. This normalization ensures that each segment of the data, organized by its periodic frequency, has a mean of zero.
A.8 Runtime
@@
In the unconditional model, we observed a neglectable computational overhead when using optimal transport maps during training. For evaluation, however, it takes approximately three times longer than the standard conditional version. We attribute this increase to two factors: the use of a longer context window (see Sec. A.2) and the need to differentiate through the ODE integration process.
A.9 Additional Forecasts
@@
A.10 Algorithms
A.10.1 Training Algorithms for TSFlow
@@
1:Input: Conditional prior distribution $q_{0}$, data distribution $q_{1}$, noise level $\sigma_{t}$, network $u_{\bm{\theta}}$
@@
6:     $\mathbf{x}_{t}\sim\mathcal{N}(\bm{\mu}_{t},\sigma_{t}^{2}\mathbf{I})$triangleright sample from $p_{t}(\cdot\mid\mathbf{z})$
7:     $\mathcal{L}({\bm{\theta}})\leftarrow\|u_{\bm{\theta}}(t,\mathbf{x}_{t},\mathbf{y}^{p})-u_{t}(\mathbf{x}_{t}\mid\mathbf{z})\|^{2}$triangleright regress conditional vector field
@@
A.11 Evaluation Algorithms for TSFlow
We provide a formal description of the conditional prior sampling algorithm in Alg. 3.
@@ A.11 Evaluation Algorithms for TSFlow
4:     $\mathbf{x}_{0}^{(i+1)}=\mathbf{x}_{0}^{(i)}-\eta\nabla_{\mathbf{x}_{0}}\log q_{0}(\mathbf{x}_{0}^{(i)}\mid\mathbf{y}^{p})+\sqrt{2\eta}\xi_{i}$triangleright update $\mathbf{x}_{0}$
@@ Appendix B Additional Results
B.1 Effect on the Optimal Transport Problem.
Our results in Fig. 6 show that Gaussian process priors reduce the Wasserstein distance between the prior and data distribution if the kernel exploits characteristics of the data. When we parametrize the prior with SE and OU kernels, the Wasserstein distance to the data distribution decreases substantially, although it slowly converges to an isotropic prior as the sequence length increases. Notably, the periodic kernel achieves the largest reduction in distance on these datasets with only marginal growth as the sequence length increases because it effectively captures the periodicity — a common feature of real-world time series — especially well. These findings support our hypothesis that an informed choice of prior distribution can substantially affect the learning problem.
B.2 Ablation
@@
B.3 Conditional Prior Sampling
@@
0 $\mathbf{0.049}$±$\pm 0.001$ $0.012$±$\pm 0.001$ $0.294$±$\pm 0.002$ $\mathbf{0.029}$±$\pm 0.001$ $\mathbf{0.371}$±$\pm 0.006$ $0.096$±$\pm 0.000$ $0.154$±$\pm 0.001$ $0.290$±$\pm 0.004$
@@
B.4 Guidance Score Function
In Eqs. 9 and 12, we utilize an asymmetric Laplace distribution (ALD) to model the guidance score. Alternatively, this score could be modeled using a Gaussian distribution, which simplifies to a mean squared error loss after applying the logarithm.
@@ B.4 Guidance Score Function
Distr. Electr. Exchange KDDCup M4Hourly Solar Traffic UberTLC Wiki2000
Gaussian $0.059$±$\pm 0.001$ $0.020$±$\pm 0.005$ $0.317$±$\pm 0.016$ $\underline{0.031}$±$\pm 0.000$ $\mathbf{0.370}$±$\pm 0.010$ $0.101$±$\pm 0.001$ $0.169$±$\pm 0.005$ $24.046$±$\pm 6.368$
ALD $\mathbf{0.049}$±$\pm 0.001$ $\mathbf{0.011}$±$\pm 0.001$ $\mathbf{0.299}$±$\pm 0.004$ $\mathbf{0.029}$±$\pm 0.001$ $0.464$±$\pm 0.004$ $\mathbf{0.089}$±$\pm 0.000$ $\mathbf{0.153}$±$\pm 0.002$ $\mathbf{0.279}$±$\pm 0.007$
@@
Table 9: Effect of the score function on the forecasting results (CRPS) for TSFlow-Uncond.  on eight real-world datasets. Best scores in bold.
As demonstrated in Tab. 9, the asymmetric Laplace distribution yields better results across 6 out of 8 datasets. We attribute this to the fact that different quantile parameters $\kappa$ take the whole distribution into account rather than focusing on a single point estimate.
B.5 Non-Gaussian Prior Distributions
In addition to Gaussian process priors, TSFlow support non-Gaussian priors for $q_{0}(\mathbf{x}^{f}_{0}\mid\mathbf{y}^{p})$. Instead of parameterizing the prior via Gaussian process regression, we define it as
$q_{0}(\mathbf{x}^{f}_{0}\mid\mathbf{y}^{p})=\mathcal{N}\left(\mathbf{x}^{f}_{0}\mid g(\mathbf{y}^{p}),\mathbf{I}\right),$
where $g$ is a base forecaster. We demonstrate this approach using the Seasonal Naive (SN) forecaster and compare it to the Ornstein-Uhlenbeck prior in Tab. 10.
Prior Electr. Exchange KDDCup M4 (H) Solar Traffic UberTLC Wiki
SN $\mathbf{0.044}$$\pm 0.001$ $\mathbf{0.008}$$\pm 0.000$ $0.321$$\pm 0.010$ $0.036$$\pm 0.001$ $0.348$$\pm 0.006$ $\mathbf{0.082}$$\pm 0.000$ $0.154$$\pm 0.002$ $0.233$$\pm 0.002$
OU $0.045$$\pm 0.001$ $0.009$$\pm 0.001$ $\mathbf{0.273}$$\pm 0.004$ $\mathbf{0.025}$$\pm 0.001$ $\mathbf{0.343}$$\pm 0.002$ $0.083$$\pm 0.000$ $\mathbf{0.153}$$\pm 0.001$ $\mathbf{0.227}$$\pm 0.000$
Table 10: Forecasting performance (CRPS) on eight datasets for the seasonal naive (SN) prior compared to the Ornstein-Uhlenbeck prior. Best scores are in bold.
B.6 Neural Function Evaluations
In Tab. 1, we present the Wasserstein-2 distances for 4 and 16 Neural Function Evaluations (NFEs). To further illustrate the performance, we show Fig. 7 the Wasserstein-2 distances across varying NFEs for selected datasets. As demonstrated in the figure, the non-isotropic priors achieve better performance than the isotropic prior.
Figure 7: Wasserstein $W_{2}^{2}$ distances between samples from unconditional generative model and data on common benchmark datasets. x-axis shows NFEs using an Euler solver.
B.7 Forecasting with missing values
In our experiments, we focus on forecasting due to the availability of standardized benchmarks and to ensure fair comparisons with autoregressive models, which do not benefit from bidirectional information passing. However, TSFlow is not restricted to a forecasting setting and can be extended to arbitrary conditional generation tasks by adjusting the observation mask.
To demonstrate this, we evaluate forecasting scenarios with missing values in the context window. Following the setup by Kollovieh et al. (2023), we test three missing values scenarios with 50% missing rate: (1) random missing (RM), where observed values are randomly masked; (2) blackout missing beginning (BO-B), where values are missing at the start of the context window; and (3) blackout missing end (BO-E), where values are missing at the end of the context window just before the forecast. For these experiments, we use the conditional version of TSFlow with the OU kernel and compare it against TSDiff, reporting the best result (conditional or unconditional) for TSDiff in each case. The results for the three setups are presented in Tabs. 11, 12 and 13, respectively.
Table 11: Forecasting results (CRPS) with missing values in the context window (RM). Best scores in bold.
Method Electricity Exchange KDDCup Solar Traffic UberTLC
TSFlow $\mathbf{0.048}$±$\pm 0.001$ $\mathbf{0.009}$±$\pm 0.001$ $0.432$±$\pm 0.121$ $0.362$±$\pm 0.006$ $\mathbf{0.085}$±$\pm 0.000$ $\mathbf{0.157}$±$\pm 0.002$
TSDiff $0.052$±$\pm 0.001$ $0.012$±$\pm 0.004$ $\mathbf{0.397}$±$\pm 0.042$ $\mathbf{0.357}$±$\pm 0.023$ $0.097$±$\pm 0.003$ $0.180$±$\pm 0.015$
Table 12: Forecasting results (CRPS) with missing values in the context window (BO-B). Best scores in bold.
Method Electricity Exchange KDDCup Solar Traffic UberTLC
TSFlow $\mathbf{0.047}$±$\pm 0.001$ $\mathbf{0.008}$±$\pm 0.001$ $\mathbf{0.305}$±$\pm 0.013$ $0.396$±$\pm 0.026$ $\mathbf{0.083}$±$\pm 0.000$ $\mathbf{0.158}$±$\pm 0.008$
TSDiff $0.049$±$\pm 0.001$ $0.009$±$\pm 0.000$ $0.441$±$\pm 0.096$ $\mathbf{0.377}$±$\pm 0.017$ $0.094$±$\pm 0.005$ $0.181$±$\pm 0.009$
Table 13: Forecasting results (CRPS) with missing values in the context window (BO-E). Best scores in bold.
Method Electricity Exchange KDDCup Solar Traffic UberTLC
TSFlow $\mathbf{0.061}$±$\pm 0.001$ $0.056$±$\pm 0.004$ $0.384$±$\pm 0.033$ $0.418$±$\pm 0.014$ $\mathbf{0.106}$±$\pm 0.002$ $\mathbf{0.165}$±$\pm 0.002$
TSDiff $0.065$±$\pm 0.003$ $\mathbf{0.020}$±$\pm 0.001$ $\mathbf{0.344}$±$\pm 0.012$ $\mathbf{0.376}$±$\pm 0.036$ $0.123$±$\pm 0.023$ $0.179$±$\pm 0.013$
Overall, TSFlow outperforms both versions of TSDiff in 12 out of 18 cases, demonstrating its robustness across these scenarios.
B.8 Multivariate Forecasting
We extend TSFlow to the multivariate setting by incorporating a transformer across the feature dimension, following the approach of CSDI (Tashiro et al., 2021). To evaluate its effectiveness, we compare this multivariate extension (MV) to a channel-independent variant (CIDP) in Table 14. For reference, we also report the forecasting performance (CRPS-Sum) of CSDI (Tashiro et al., 2021) and the method proposed by Biloš et al. (2023).
Model Electricity Exchange Solar Traffic
Biloš et al. (2023) $0.0260$$\pm 0.0011$ $0.0071$$\pm 0.0009$ $0.3701$$\pm 0.0179$ $1.2679$$\pm 1.0624$
CSDI $0.0184$$\pm 0.0008$ $0.0076$$\pm 0.0022$ $0.3179$$\pm 0.0197$ $0.0221$$\pm 0.0012$
TSFlow (CIDP) $0.0196$$\pm 0.0006$ $\mathbf{0.0060}$$\pm 0.0004$ $0.3780$$\pm 0.0074$ $0.0240$$\pm 0.0005$
TSFlow (MV) $\mathbf{0.0176}$$\pm 0.0012$ $0.0073$$\pm 0.0005$ $\mathbf{0.2792}$$\pm 0.0154$ $\mathbf{0.0217}$$\pm 0.0013$
Table 14: Forecasting performance (CRPS-Sum) on four datasets. Best scores are in bold.
The results show that yields the best results on 3 out of 4 datasets.
@@ Appendix C Guided Generation
To enable guided generation, we use slightly modified conditional probability paths defined as:
$p_{t}(\mathbf{x}_{t}\mid\mathbf{z})=\mathcal{N}\left(\mathbf{x}_{t};\bm{\mu}_{t},\sigma_{t}^{2}\mathbf{I}\right),$ (19)
where $\bm{\mu}_{t}=(1-t)\mathbf{x}_{0}+t\mathbf{x}_{1}$, $\sigma_{t}=(1-t)\sigma_{\mathrm{max}}+t\sigma_{\mathrm{min}}$, and $\sigma_{\mathrm{max}}$ and $\sigma_{\mathrm{min}}$ are constant standard deviations with $\sigma_{\mathrm{max}}>\sigma_{\mathrm{min}}$. Note that when $\sigma_{\mathrm{max}}=\sigma_{\mathrm{min}}$, we recover the same probability path as in Sec. 2.2. However, we require $\sigma_{\mathrm{max}}>\sigma_{\mathrm{min}}$ to obtain a score function in our vector field.
@@
$u_{t}(\mathbf{x}_{t}\mid\mathbf{z})=\mathbf{x}_{1}-\mathbf{x}_{0}+\frac{\sigma_{\mathrm{min}}-\sigma_{\mathrm{max}}}{(1-t)\sigma_{\mathrm{max}}+t\sigma_{\mathrm{min}}}\left(\mathbf{x}_{t}-\mu_{t}\right).$ (20)
@@
$u_{t}(\mathbf{x}_{t}\mid\mathbf{z})=\mathbf{x}_{1}-\mathbf{x}_{0}-(\sigma_{\mathrm{min}}-\sigma_{\mathrm{max}})\sigma_{t}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}\mid\mathbf{z})$ (21)
@@
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}\mid\mathbf{z})=-\frac{\mathbf{x}_{t}-(1-t)\mathbf{x}_{0}-t\mathbf{x}_{1}}{\left((1-t)\sigma_{\mathrm{max}}+t\sigma_{\mathrm{min}}\right)^{2}}.$ (22)
@@
$=$ $\int\frac{(\mathbf{x}_{1}-\mathbf{x}_{0}-(\sigma_{\mathrm{min}}-\sigma_{\mathrm{max}})\sigma_{t}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}\mid\mathbf{z}))p_{t}(\mathbf{x}_{t}\mid\mathbf{z})q(\mathbf{z})}{p_{t}(\mathbf{x}_{t})}\mathrm{d}\mathbf{z}$ (24)
$=$ $\int\frac{(\mathbf{x}_{1}-\mathbf{x}_{0})p_{t}(\mathbf{x}_{t}\mid\mathbf{z})q(\mathbf{z})}{p_{t}(\mathbf{x}_{t})}\mathrm{d}\mathbf{z}$ (25)
-- $(\sigma_{\mathrm{min}}-\sigma_{\mathrm{max}})\sigma_{t}\int\frac{\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}\mid\mathbf{z})p_{t}(\mathbf{x}_{t}\mid\mathbf{z})q(\mathbf{z})}{p_{t}(\mathbf{x}_{t})}\mathrm{d}\mathbf{z}$ (26)
$=$ $v_{t}(\mathbf{x}_{t})-(\sigma_{\mathrm{min}}-\sigma_{\mathrm{max}})\sigma_{t}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}),$ (27)
@@
$\tilde{u}_{t}(\mathbf{x}_{t})=\underbrace{v_{t}(\mathbf{x}_{t})-(\sigma_{\mathrm{min}}-\sigma_{\mathrm{max}})\sigma_{t}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})}_{=u_{t}(x_{t})}-s(\sigma_{\mathrm{min}}-\sigma_{\mathrm{max}})\sigma_{t}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}^{p}\mid\mathbf{x}_{t}),$ (29)
where $s$ is a scaling parameter introduced to control the strength of the conditioning. We can summarize this as:
$\tilde{u}_{t}(\mathbf{x}_{t})=u_{t}(\mathbf{x}_{t})-s\sigma_{t}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}^{p}\mid\mathbf{x}_{t}),$ (30)
[…]